# Explaining word embeddings with perfect fidelity: Case study in research impact prediction

### v1

@@ Abstract
Best performing approaches for scholarly document quality prediction are based on embedding models, which do not allow direct explanation of classifiers as distinct words no longer correspond to the input features for model training. Although model-agnostic explanation methods such as Local interpretable model-agnostic explanations (LIME) can be applied, these produce results with questionable correspondence to the ML model. We introduce a new feature importance method, Self-model Rated Entities (SMER), for logistic regression-based classification models trained on word embeddings. We show that SMER has theoretically perfect fidelity with the explained model, as its prediction corresponds exactly to the average of predictions for individual words in the text. SMER allows us to reliably determine which words or entities positively contribute to predicting impactful articles. Quantitative and qualitative evaluation is performed through five diverse experiments conducted on 50.000 research papers from the CORD-19 corpus. Through an AOPC curve analysis, we experimentally demonstrate that SMER produces better explanations than LIME for logistic regression.
@@ Introduction
Word embeddings perform well in tasks related to the identification of impactful research. This included the task of citation prediction [4] as well as predicting whether a paper will be accepted or rejected [3, 2]. However, the evaluations of these approaches have focused largely on model predictive performance, while many applications require reliable explanations. Having the information that a draft of the paper will not be accepted is not as useful to the author as a specific and reliable explanation of what specifically contributes to its positive or negative perception by the model. We investigated the use of several state-of-the-art model-agnostic methods for explanations, including LIME and Shapley plots, in our previous work [5], noting that the quality of the explanations generated for why specific research was cited lags behind white-box models such as rule learning. The disadvantage of these methods is that they obscure the explanation process by introducing randomly generated data and additional machine-learning models. The quality of such explanations is not yet thoroughly studied, with research showing that methods such as LIME can be manipulated to have very low fidelity [6].
In this article, we, therefore, propose a new explanation method that provides intrinsic explanations of embedding-based predictive models without training additional models or generating new instances. Furthermore, we theoretically show that the newly proposed explanation method is guaranteed to have perfect fidelity for logistic regression predictive models. In the five experiments, we confirm this result empirically.
@@
This section describes the workflow of our methodology for predicting the impactfulness of words as well as entire documents. A reliable prediction of research impact for any research paper only based on its abstract is a difficult task. In our research, we address a narrower subproblem of predicting which research will be groundbreaking, ie., we are interested in identifying the top most influential research – both articles and newly emerged entities, such as drugs or virus strains in the CORD-19 training data context. The focus on the most confident predictions makes the research impact and explainability problem more tractable.
We first describe our new method of building an embedding-based classifier for predicting article impact, which will directly provide word-based explanations. Then, we will cover two existing baseline methods (ORC and LIME). Based on this, we will provide an overview of the framework for predicting the impact of research articles and impactful entities (words). Finally, we will cover the methods we used for evaluation in the five experiments.
@@
We represent a scientific article by its abstract, which is a string consisting of dictionary words. Let $\mathcal{W}$ stand for a dictionary; an element $w\in\mathcal{W}$ is called a word. A document, a text—or specifically in this study, an abstract of a scientific paper—is a nonempty sequence $A=(w_{1},\dots,w_{N})$ of words. We are given a family $\mathcal{A}$ of abstracts $A_{i}=(w_{i,1},\dots,w_{i,N_{i}})$, $i=1,\dots,n$, called corpus. We assume that $\mathcal{W}=\bigcup_{i=1}^{n}\bigcup_{j=1}^{N_{i}}w_{i,j}$.
@@
The most significant property of embeddings for this study is that when words $w,w^{\prime}\in\mathcal{W}$ frequently appear “close” to each other in the training dataset $\mathcal{A}$, then $D(\widetilde{w},\widetilde{w}^{\prime})$ is “small”, where $D$ stands for some distance measure on $\mathbb{R}^{d}$, such as the cosine similarity or Euclidian distance.
@@
Recall that an abstract $A$ of a scientific paper is here formalized simply as a sequence of some dictionary words $w_{1},\dots,w_{N}$, possibly with repetitions. The aim is to construct $\widetilde{A}\in\mathbb{R}^{d}$, an embedding of $A$, a representation of abstract as a vector in $\mathbb{R}^{d}$. The function $A\mapsto\widetilde{A}$ should fulfil certain natural properties, such as
  1. 1.
    (reflexivity) if $A=(w,w,\dots,w)$, then $\widetilde{A}=\widetilde{w}$;
  2. 2.
    (replacement principle) if $A=(w_{1},\dots,w_{j-1},w_{j},w_{j+1},\dots,w_{N})$,
    $A^{\prime}=(w_{1},\dots,w_{j-1},w_{j}^{\prime},$ $w_{j+1},\dots,w_{N})$ and $\|\widetilde{w}_{j}-\widetilde{w}_{j}^{\prime}\|\leq\varepsilon$, then $\|\widetilde{A}-\widetilde{A}^{\prime}\|\leq\varepsilon/N$;
  3. 3.
    (extension principle) if $A=(w_{1},\dots,w_{j},w_{j+1},\dots,w_{N})$, $A^{\prime}=(w_{1},\dots,w_{j},\\ w^{\prime},w_{j+1},\dots,w_{N})$ and $\|w^{\prime}-w_{\iota}\|\leq\varepsilon$ for every ${\iota}=1,\dots,N$, then $\|\widetilde{A}-\widetilde{A}^{\prime}\|\leq\varepsilon/N$.
Reflexivity states that if an abstract $A$ contains a single dictionary word (possibly with multiple occurrences), then the embedding of that abstract coincides with the embedding of the word. The replacement principle is a form of continuity: if we replace a word $w_{j}$ in an abstract with a close word $w_{j}^{\prime}$, then the embedding of the abstract cannot change too much. Finally, the extension principle is another form of continuity: if we have an abstract $A$ and we extend it by adding a new word $w^{\prime}$, which is close to all words $w_{1},\dots,w_{N}$, then the embedding of the extended abstract $A^{\prime}$ cannot be too far from $A$.
In this study, we aggregate embeddings of words appearing in an abstract as follows: if $A=(w_{1},\dots,w_{N})$, then
$\widetilde{A}=\frac{1}{N}\sum_{j=1}^{N}\widetilde{w}_{j},$ (1)
which we refer to as weighted-average embedding. (The average is weighted in the sense that a word is counted in the above sum as many times as it appears in $A$.) For dataset $\mathcal{A}$, we have $\widetilde{A}_{i}=N_{i}^{-1}\sum_{j=1}^{N_{i}}\widetilde{w}_{i,j}$, $i=1,\dots,n$.
Remark
It is easy to see that (Eq. 1) fulfills properties 1–3. Reflexivity is obvious. The replacement principle follows from the fact that $\|\widetilde{A}-\widetilde{A}^{\prime}\|=\|\frac{1}{N}\sum_{\iota=1}^{N}\widetilde{w}_{\iota}-\frac{1}{N}[\widetilde{w}_{j}^{\prime}+\sum_{\iota\in\{1,\dots,N\}\setminus\{j\}}\widetilde{w}_{\iota}]\|=\|\frac{1}{N}(\widetilde{w}_{j}-\widetilde{w}_{j}^{\prime})\|\leq\varepsilon/N$. The extension principle is proved analogously: $\|\widetilde{A}-\widetilde{A}^{\prime}\|=\|\frac{1}{N}\sum_{\iota=1}^{N}\widetilde{w}_{\iota}-\frac{1}{N+1}[\widetilde{w}^{\prime}+\sum_{\iota=1}^{N}\widetilde{w}_{\iota}]\|=\|\frac{1}{N(N+1)}\sum_{\iota=1}^{N}\widetilde{w}_{\iota}-\frac{N}{N(N+1)}\widetilde{w}^{\prime}\|=\|\frac{1}{N(N+1)}\sum_{\iota=1}^{N}(\widetilde{w}_{\iota}-\widetilde{w}^{\prime})\|\leq\|\frac{1}{N(N+1)}N\varepsilon\|=\varepsilon/(N+1)\leq\varepsilon/N.$
@@
$\mathsf{E}[y_{i}\mid A_{i}]=L(\beta_{0}+\beta^{T}\widetilde{A}_{i}),\quad i=1,\dots,n,$
@@
The basis of our self-explanatory embeddings-based classification framework is the observation that a word $w\in\mathcal{W}$ can be seen as a single-word abstract $A=(w)$. We can thus compute feature importance (i.e. word importance) directly using the same model, which was trained to classify all documents (the SCORE from Eq. 3). When applied to a single word, the formula reduces to
@@
Words $w\in\mathcal{W}$ can be sorted according to their scores; by selection of a cutoff threshold, we get a classifier of high- and low-scoring words, i.e., words that have high or low potential to contribute to the number of article’s citations. Since $\widetilde{w}$ is an embedding vector, the method can also score words unseen in the training data as long as these can be assigned an embedding vector.
@@
Proposition 1 is a key property for assessing the fidelity of SMER.
Fidelity is defined as the percentage of predictions that are the same between the original model and the explanation [8, 9]. Proposition 1 tells us that the average of the individual explanations (predictions for word vectors of individual words) is equal to the prediction of the explained model (the document classifier). As a consequence, the SMER explanation technique has 100% fidelity.
Proposition 2 (preservation of ordering)
@@
$\textit{SCORE}(A)$ $=L(w(A))=L\left(\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell})\right)<L\left(\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell})+\frac{\Delta}{N}\right)$
$=L\left(\frac{1}{N}\left(\sum_{\ell\in\{1,\dots,N\}\setminus\{k\}}\lambda(w_{\ell})\right)+\frac{\lambda(w^{\prime}_{k})}{N}\right)=L(w(A^{\prime}))=\textit{SCORE}(A^{\prime}).$
@@
Observe that $\widehat{\gamma}_{k}>0$ indicates that adding the word $v_{k}$ into an abstract (if not yet present) increases the odds of the article becoming highly cited. On the other hand, $\widehat{\gamma}_{k}<0$ indicates that adding $v_{k}$ into an abstract is predicted to decrease the odds of the article becoming highly cited. Therefore, the coefficient $\widehat{\gamma}_{k}$ measures the impact of the dictionary word $v_{k}$ on the target variable $y$. Recall that within the logistic regression framework this effect can be interpreted in terms of odds ratio. We define
@@
Figure 1: Methodological pipeline of workflow steps.
@@
The data utilized for this study is derived from the CORD-19 corpus [11], which consists primarily of research articles pertaining to biology and medicine relevant to SARS-CoV-2 and that were published between 1952-2023. Specifically, we employed the version of the corpus downloaded on April 23, 2022, comprising a total of 992,921 articles. Article citation counts were extracted from Opencitations.org.
For the training of classifiers, we used only article abstracts from the years 2016-2021 for which the OpenCitations data were available. The data sizes used for individual experiments are summarized in Table 1. As the table shows, more article abstracts were available for the training of embeddings because, for this task, labels derived from OpenCitations were not required.
Publication years Used in Count of abstracts for training of
Word embeddings Classification
2016+2017 Experiment 1 12,589 1,867
2017+2018 Experiment 1 13,036 1,916
2018+2019 Experiment 1 14,503 2,256
2019+2020 Experiment 1 + 2 +3 132,422 18,571
2020+2021 Experiment 4 + 5 414,174 6,964
Table 1: Number of abstracts used in training embedding models and classification models.
@@
The number of articles was reduced to 598,587 after filtering out abstracts with duplicate document object identifiers (DOI). Subsequently, articles whose number of characters did not exceed 10 were removed, reducing the article count to 483,781 articles. Filtering of non-English abstracts detected using the pre-trained Fasttext [12, 13, 14] language identification model reduced the count to 477,237 articles. For training, only articles with recent years of publication (2016-2021) were used, which can be seen in Table 1. For training classification models, 80% of all articles with citation data for the specified years of publications were used, which is the number shown in the last column of Table 1. The total number of unique abstracts we used for training embeddings was 225,679. Note that to train word embedding models, more articles were available as abstracts with missing citation data could also be used.
The abstracts were preprocessed as follows. Punctuation was removed except for dashes (as these are used in compound biomedical names), all text was converted to lowercase, and single numbers were deleted. Lemmatization was used to convert the words into root words. After this preprocessing, articles with less than 20 characters were removed. The final dataset contained 476,175 articles. After preprocessing, we also checked for possible duplications, determining that more than 99.8% of abstracts were unique.
@@
The counts of articles used to train the embeddings and the classification models differed. For the unsupervised task of embedding training, articles without available citation counts could be used, while for learning the classifier, the availability of citation counts was required to construct the target. The sizes of the respective datasets are summarized in Table 1.
@@
After the preprocessing steps, article abstracts were represented using two variant methods: first, embeddings-based word vectors generated from abstracts with Eq.1, and second, we used an intrinsically explainable model based on Bag of Words (see Eq. 5).
However, unlike several recent approaches, we deliberately excluded the author information and other metadata, such as journal impact factor, from training since we have found in our previous work that it can lead to publication biases [5].
The details of data used to train the embedding models for individual experiments are presented in Table 1. The dimension $d\geq 1$ is a parameter used to train the embedding models. For training the word embeddings, we used embedding vector dimensions of 100 and a window size of 10. These settings were based on our prior experience combined with recommendations from the literature [16].
@@
This shows that while citations often fail to reflect the true impact for various reasons, this often happens on an individual level, which is not relevant to our work as our machine learning model learns an abstraction. Our research is based on this widely held assumption, well summarized by [18]: ”Although most citation analysts seem to agree that citing or referencing is biased, it has been argued that this bias is not fatal for the use of citation as performance indicators—to a certain extent, the biases are averaged out at aggregated levels … the presence of different cognitive meanings of citations and motivations for citing does not necessarily invalidate the use of citations as (imperfect) performance measures…”.
@@
To define the target, we combine two subtasks of scholarly document quality prediction: predicting citation counts and predicting whether a paper will be rejected or accepted [3]. We define the target variable $y_{i}$ as a 0-1 indicator of being a highly cited article. To define the notion precisely, let
$\textit{CIT}(i,t)$ $:=\text{number of citations of article $A_{i}$ originated in year$t$},$ (7)
$\textit{PUB}(i)$ $:=\text{year of publication of article $A_{i}$.}$ (8)
@@
Recall that $n$ stands for the number of corpus articles. Both training and evaluation are performed on distinct time slices defined by the set $T$ of publications years, e.g. $T=\{2016,2017\}$ and a fixed citation year $\tau$, e.g. $\tau=2018$. Then, article $A_{i}$ with $\textit{PUB}(i)\in T$ is defined to be highly cited with respect to $T$ and $\tau$ if the number of its citations from year $\tau$ (strictly) exceeds the median:
@@
Table 2 presents the composition of the low- and highly-cited articles in the dataset after the target has been derived. The count of articles that were published in a given year differs in individual citation years due to missing citation data. These were not consistently available for each article in all covered years. Some articles are cited before they have been published — this situation can occur at the preprint stage of the article. Following the strict inequality in definition (12), the two subsets separated by the median are not completely equally sized. For example, assume a corpus containing five articles with citation counts $0,1,1,1,2$, then as a result of the preprocessing, there will be only one article in the “high” category, i.e. article with the citation count strictly above the median $1$, and the remaining four articles with at most one citation will be assigned to the “low” category.
@@
Table 2: Number of low- and highly-cited articles by publication and citation year in the used subset of the CORD-19 corpus used for training and evaluation of classification models.
Evaluation methodology
Data-splitting strategies
@@
The time-based nature of the data requires a specific approach to data splitting. We report results using the time-split and random-split techniques, two commonly used approaches. Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest articles, which negatively affects their ability to evaluate the impact of recent research. We, therefore, also report results using a third technique, which addresses this issue, which we call semi-split testing.

### v2

# Explaining word embeddings with perfect fidelity: A case study in predicting research impact

@@ Abstract
The best-performing approaches for scholarly document quality prediction are based on embedding models. In addition to their performance when used in classifiers, embedding models can also provide predictions even for words that were not contained in the labelled training data for the classification model, which is important in the context of the ever-evolving research terminology. Although model-agnostic explanation methods, such as Local interpretable model-agnostic explanations, can be applied to explain machine learning classifiers trained on embedding models, these produce results with questionable correspondence to the model. We introduce a new feature importance method, Self-model Rated Entities (SMER), for logistic regression-based classification models trained on word embeddings. We show that SMER has theoretically perfect fidelity with the explained model, as the average of logits of SMER scores for individual words (SMER explanation) exactly corresponds to the logit of the prediction of the explained model. Quantitative and qualitative evaluation is performed through five diverse experiments conducted on 50,000 research articles (papers) from the CORD-19 corpus. Through an AOPC curve analysis, we experimentally demonstrate that SMER produces better explanations than LIME, SHAP and global tree surrogates.
@@ Introduction
Word embeddings perform well in tasks related to the identification of impactful research. This included the task of citation prediction [4] as well as predicting whether a paper will be accepted or rejected [3, 2]. However, the evaluations in these approaches have focused largely on model predictive performance, while many applications require reliable explanations. Having the information that a draft of the paper will not be accepted is not as useful to the author as a specific and reliable explanation of what specifically contributes to its positive or negative perception by the model. We previously investigated several state-of-the-art model-agnostic methods for explanations, including LIME and Shapley plots, in our previous work [5], noting that the quality of the explanations generated for why specific research was cited lags behind white-box models such as rule learning. These methods often obscure the explanation process by introducing randomly generated data and additional machine-learning models. The quality of such explanations is not yet thoroughly studied, with research showing that methods such as LIME can be manipulated to have very low fidelity [6].
In this publication, we, therefore, propose a new explanation method that provides intrinsic explanations of embedding-based predictive models without training additional models or generating new instances. Furthermore, we theoretically show that the newly proposed explanation method is guaranteed to have perfect fidelity for logistic regression predictive models. We empirically confirm this result in five experiments.
@@
This section describes the workflow of our methodology for predicting the impactfulness of words as well as entire documents. A reliable prediction of research impact for any research paper only based on its abstract is a difficult task. In our research, we address a narrower subproblem of predicting which research will be groundbreaking, that is, we are interested in identifying the most influential research – both papers and newly emerged entities, such as drugs or virus strains in the CORD-19 training data context. The focus on the most confident predictions makes the research impact and explainability problem more tractable.
We first describe our new method of building an embedding-based classifier for predicting paper impact, which will directly provide word-based explanations. Then, we will cover two existing baseline methods (ORC and LIME). Based on this, we will provide an overview of the framework for predicting the impact of research papers and impactful entities (words). Finally, we will cover the methods we used for evaluation in the five experiments.
@@
We represent a scientific paper by its abstract, which is a string consisting of dictionary words. Let $\mathcal{W}$ stand for a dictionary; an element $w\in\mathcal{W}$ is called a word. A document, a text—or specifically in this study, an abstract of a scientific paper—is a nonempty sequence $A=(w_{1},\dots,w_{N})$ of words. We are given a family $\mathcal{A}$ of abstracts $A_{i}=(w_{i,1},\dots,w_{i,N_{i}})$, $i=1,\dots,n$, called corpus. We assume that $\mathcal{W}=\bigcup_{i=1}^{n}\bigcup_{j=1}^{N_{i}}w_{i,j}$.
@@
The most significant property of embeddings for this study is that when words $w,w^{\prime}\in\mathcal{W}$ frequently appear “close” to each other in the training dataset $\mathcal{A}$, then $D(\widetilde{w},\widetilde{w}^{\prime})$ is “small”, where $D$ stands for some distance measure on $\mathbb{R}^{d}$, such as the cosine similarity or Euclidean distance.
@@
Having constructed an embedding $w\in\mathcal{W}\mapsto\widetilde{w}\in\mathbb{R}^{d}$ of words, the next step is a construction of an embedding of abstracts, which is a function $A\mapsto\widetilde{A}\in\mathbb{R}^{d}$, where $A$ is an abstract — a sequence of words $A=(w_{1},\dots,w_{N}).$ We use the average embedding
$\widetilde{A}=\frac{1}{N}\sum_{j=1}^{N}\widetilde{w}_{j}.$ (1)
This aggregation function has the following favorable property of linearity: $\widetilde{A}$ is the average of the vectors $\widetilde{w}_{1},\dots,\widetilde{w}_{N}$, a feature essential for the proofs of Propositions 1–3.
@@
$\mathsf{E}[y_{i}\mid A_{i}]=L(\beta_{0}+\beta^{T}\widetilde{A}_{i}),\quad i=1,\dots,n,$ (2)
@@
The basis of our self-explanatory embedding-based classification framework is the observation that a word $w\in\mathcal{W}$ can be seen as a single-word abstract $A=(w)$. We can thus compute feature importance (i.e. word importance) directly using the same model, which was trained to classify all documents (the SCORE from Eq. 4). When applied to a single word, the formula reduces to
@@
Words $w\in\mathcal{W}$ can be sorted according to their scores; by selection of a cutoff threshold, we get a classifier of high- and low-scoring words, i.e., words that have high or low potential to contribute to the number of paper’s citations. Since $\widetilde{w}$ is an embedding vector, the method can also score words unseen in the training data as long as these can be assigned an embedding vector.
@@
Proposition 1 is a key property for assessing the fidelity.
Fidelity is the percentage of predictions that are the same between the original model and the explanation [8, 9]. Formally, for a finite evaluation set $\mathcal{D}\subset\mathcal{A}$:
$\mathrm{Fidelity}(g,f;\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{A\in\mathcal{D}}\mathbf{1}\bigl{[}g(A)=f(A)\bigr{]}.$
Here, $f\colon\mathcal{A}\to\mathbb{R}$ is the prediction of the original model, and $g\colon\mathcal{A}\to\mathbb{R}$ is the prediction of the explanation method being evaluated.
In our case, fidelity holds in the logit space, where
$f(A)=\lambda(A),\qquad g(A)=\tfrac{1}{N}\sum_{j=1}^{N}\lambda(w_{j}),$
and Proposition 1 ensures
$f(A)=g(A)\quad\text{for all }A\in\mathcal{D}.$
For presentation, we apply the logistic function $L(\cdot)$:
$\textit{SCORE}(A):=L(f(A)),\qquad\textit{SCORE}(w_{j}):=L(\lambda(w_{j})).$
This rescaling improves interpretability, and since $L$ is strictly increasing (Proposition 2), the ranking of words is preserved.
Proposition 2 (order preservation)
@@
$\textit{SCORE}(A)$ $=L\bigl{(}\lambda(A)\bigr{)}=L\left(\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell})\right)<L\left(\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell})+\frac{\Delta}{N}\right)$
$=L\left(\frac{1}{N}\!\left(\sum_{\ell\in\{1,\dots,N\}\setminus\{k\}}\lambda(w_{\ell})\right)+\frac{\lambda(w^{\prime}_{k})}{N}\right)=L\bigl{(}\lambda(A^{\prime})\bigr{)}=\textit{SCORE}(A^{\prime}).$
@@
Observe that $\widehat{\gamma}_{k}>0$ indicates that adding the word $v_{k}$ into an abstract (if not yet present) increases the odds of the paper to become highly cited. On the other hand, $\widehat{\gamma}_{k}<0$ indicates that adding $v_{k}$ into an abstract is predicted to decrease the odds of the paper becoming highly cited. Therefore, the coefficient $\widehat{\gamma}_{k}$ measures the impact of the dictionary word $v_{k}$ on the target variable $y$. Recall that within the logistic regression framework this effect can be interpreted in terms of odds ratio. We define
@@
Figure 1: The figure shows the methodological pipeline of workflow steps.
@@
The data utilized for this study is derived from the CORD-19 corpus [11], which consists primarily of research papers pertaining to biology and medicine relevant to SARS-CoV-2 and that were published between 1952-2023. Specifically, we employed the version of the corpus downloaded on April 23, 2022, comprising a total of 992,921 papers. Paper citation counts were extracted from Opencitations.org. For the training of classifiers, we used only paper abstracts from the years 2016-2021 for which the Opencitation data were available.
@@
The number of papers was reduced to 598,587 after filtering out abstracts with duplicate document object identifiers (doi). Subsequently, papers whose number of characters did not exceed 10 were removed, reducing the paper count to 483,781 papers. Filtering of non-English abstracts detected using the pre-trained Fasttext [12, 13, 14] language identification model reduced the count to 477,237 papers. For training embedding models, only papers with recent years of publication (2016-2021) were used and for training classification models, 80% of all papers with citation data for the specified years of publications were used. The total number of unique abstracts we used for training embeddings was 225,679. Note that to train word embedding models, more papers were available as abstracts with missing citation data could also be used.
The abstracts were preprocessed as follows. Punctuation was removed except for dashes (as these are used in compound biomedical names), all text was converted to lowercase, and single numbers were deleted. Lemmatization was used to convert the words into root words. After this preprocessing, papers with less than 20 characters were removed. The final dataset contained 476,175 papers. After preprocessing, we also checked for possible duplications, determining that more than 99.8% of abstracts were unique.
@@
After the preprocessing steps, paper abstracts were represented using two variant methods: first, embedding-based word vectors generated from abstracts with Eq. 1, and second, we used an intrinsically explainable model based on bag-of-words (BoW) (see Eq. 5).
However, unlike several recent approaches, we deliberately excluded the author information and other metadata such as journal impact factor, from training, since we have found in our earlier work that it can lead to publication biases [5].
The dimension $d\geq 1$ is a parameter used to train the embedding models. For training the word embeddings, we used embedding vector dimensions of 100 and a window size of 10. These settings were based on our prior experience combined with recommendations from the literature [16].
@@
This shows that while citations often fail to reflect the true impact for various reasons, this often happens on an individual level, which is not relevant to our work as our machine learning model learns an abstraction. Our research is based on this widely held assumption, well summarized by [18]: ”Although most citation analysts seem to agree that citing or referencing is biased, it has been argued that this bias is not fatal for the use of citation as performance indicators—to a certain extent, the biases are averaged out at aggregated levels … the presence of different cognitive meanings of citations and motivations for citing does not necessarily invalidate the use of citations as (imperfect) performance measures…”
@@
To define the target, we combine two subtasks of scholarly document quality prediction: predicting citation counts and predicting whether a paper will be rejected or accepted [3]. We define the target variable $y_{i}$ as a 0-1 indicator of being a highly cited paper. To define the notion precisely, let
$\textit{CIT}(i,t)$ $:=\text{number of citations of paper $A_{i}$ originated in year $t$},$ (7)
$\textit{PUB}(i)$ $:=\text{year of publication of paper $A_{i}$.}$ (8)
@@
Recall that $n$ stands for the number of corpus papers. Both training and evaluation are performed on distinct time slices defined by the set $T$ of publications years, e.g. $T=\{2016,2017\}$ and a fixed citation year $\tau$, e.g. $\tau=2018$. Then, paper $A_{i}$ with $\textit{PUB}(i)\in T$ is defined to be highly cited with respect to $T$ and $\tau$ if the number of its citations from year $\tau$ (strictly) exceeds the median:
@@
Table 1 presents the composition of the low- and highly cited papers in the dataset after the target has been derived. The count of papers that were published in a given year differs in individual citation years due to missing citation data. These were not consistently available for each paper in all covered years. Some papers are cited before they have been published — this situation can occur at the preprint stage of the paper. Following the strict inequality in definition (12), the two subsets separated by the median are not completely equally sized. For example, assume a corpus containing five papers with citation counts $0,1,1,1,2$, then as a result of the preprocessing, there will be only one paper in the “high” category, i.e. paper with the citation count strictly above the median $1$, and the remaining four papers with at most one citation will be assigned to the “low” category.
@@
Table 1: Number of low- and highly cited papers by publication and citation year in the used subset of the CORD-19 corpus used for training and evaluation of classification models.
Justification of Binary Target Selection
We also explored a regression-based approach, predicting the exact number of citations and applying a threshold to derive class labels. However, as demonstrated in our previous work [5], this approach resulted in lower predictive performance compared to the direct binary classification setup. Consequently, we adopted a classification formulation (with a binary instead of continuous target). It should be emphasised that even with a binary target (high vs low cited categories), it is still possible to discern the most significant research, since this is reflected in the probability of the classification. An article predicted 99% to be in the high category is considered more significant than an article with 51% predicted probability. In this way, the binary target formulation provides both more accurate and interpretable means for identifying impactful research compared to predicting the number of citations directly.
Experiment 1: Predictions of impactful papers
The goal of this experiment is to evaluate the predictive performance on the task of identifying the most (and least) promising research. We compare the proposed method, logistic regression over word embeddings (further referred to as Word2vec or w2v), with an older approach based on BoW.
Datasets
For models using embeddings, for each of these datasets, embeddings were trained separately on abstracts published in the corresponding years. This experiment evaluates four consecutive two-year windows (2016–2017 to 2019–2020). Each window supplies all abstracts for training word embeddings, see number of abstracts in Table 2
Publication window Word embeddings Classification
2016–2017 12 589 1 867
2017–2018 13 036 1 916
2018–2019 14 503 2 256
2019–2020 132 422 18 571
Table 2: Number of abstracts used in Experiment 1.
To train the classification models, we created four datasets, each covering papers published in two consecutive years from the period 2016-2020. The overview of the datasets is given in Table 3.
Evaluation strategies
@@
The time-based nature of the data requires a specific approach to data splitting. We report results using the time-split and random-split techniques, two commonly used approaches. Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest papers, which negatively affects their ability to evaluate the impact of recent research. We therefore also report results using a third technique, which addresses this issue, which we call semi-split testing.
Random-split testing.
Random, time-sensitive splits of data are a common evaluation technique. In our setup, 80% of the data from selected periods is used to train the models and the remaining 20% from the same periods is used for testing.
Time-split testing.
This split is stricter than random-split testing, as the testing set consists of papers that were both published and cited later than the training set in order to prevent any leakage from the test data to the training data.
For training, we use 80% of available papers from publication years $t-1,t-2,\dots$ with the target being derived from citation counts of these papers accumulated by year $t-1$. The test papers were those published in year $t$ with the target being derived from citation counts also from year $t$. This is a particularly challenging setup because the prediction of citations for papers published in the same year is often very difficult.
Semi-split testing.
Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest papers, which negatively affects their ability to evaluate the impact of recent research. This semi-split evaluation technique best corresponds to how we envisage that the models would be used by leveraging the latest available citation counts for testing. Let $t$ be the most recent publication year. For training, we take 80% papers from $t-1$ to $t$, using citation counts accumulated up to the end of year $t$. The remaining 20% of the $t$-cohort form the test set and are evaluated by the citations they collect during the following calendar year $t+1$. Thus, the model learns from the freshest vocabulary while still being judged on genuine one-year-ahead citation predictions.
Focus on groundbreaking research
To evaluate the ability of the model to separate groundbreaking research from uninfluential papers, we construct a special testing set Extreme10%. Using citation count as a proxy for the impact of the research. This is a subset of the respective test set containing a selection of 5% of papers with the highest citation counts and 5% of papers with the lowest citation counts.
Evaluation measures for the prediction task
Evaluation scenarios
The last column of Table 3 identifies the given combination of publication and citation year used in the train and test data and type of split as the evaluation scenario. For example, for training the classification model in scenario 1 (publication years 2016 and 2017), 80% of all papers published in 2016+2017 with citations in 2017 were used, which makes a total of (962+375+629+368) = 2334 x 0.8 = 1867 papers, which also corresponds to the value in Table 2.
Results
The results are presented in Table 3.